OpenStudy (anonymous):

Find the center and vertices of the hyperbola. 11x^2 - 25y^2 - 110x + 150y - 225 = 0

OpenStudy (anonymous):

Organize first... so \[11x ^{2} -110x -25y ^{2} +150y = 225\]

OpenStudy (anonymous):

\[\frac{ (x-h)^{2} }{ a ^{2} } - \frac{ (y-k)^{2} }{ b ^{2} } = 1\] can you get it into this form by completing the square for x and y.. Than divide both sides by the square..

OpenStudy (anonymous):

Alright so , did you get to the part 11(x^2 - 10x +25) - 25(y^2 +6y +9) = 225 + 25 + 9